# Mitigating Dynamic Tip-Over during Mobile Crane Slewing using Input Shaping

### v1

@@ 1 Introduction
Mobile cranes are widely utilized in construction, shipping, and industrial operations, where both safety and productivity are crucial. Despite their versatility, mobile cranes remain highly susceptible to instability, with tip-over accidents continuing to be a major cause of fatalities and equipment damage (Tip_Over_Stability_Analysis; tipover_stability_analysis_mobile; noish_alert; OSHA_MobileCraneAccidents; tipover_responses; stability_analysis). A significant contributor to such incidents is the dynamic interaction between the boom and the suspended payload during slewing motion. In practice, load limits for crane safety are prescribed by load charts that are based on static equilibrium analysis (GroveGMK7550; mobile_crane_inspector) to specify the maximum allowable load as a function of crane configuration. These limits do not account for dynamic effects such as payload swing due to slewing. As a result, crane operations that are theoretically within the static safe limits may still become unstable due to dynamic effects, as shown in Fig. 1. The goal of this work is to avoid tip-over due to excessive dynamics-induced payload swing during slewing, even with statically safe loads.
@@
Tip-over can be mitigated by modifying the crane’s mechanical configuration or by automating its motion. For example, increasing the counterweight mass has been shown to expand the stability margin and improve lifting capacity in crawler cranes (lifting_path_planning). Other approaches focus on controlling outrigger forces to stabilize the crane during motion (synchronous_control_theory), though moving outriggers during an operation is generally unsafe. Additional studies examine specific crane architectures and derive model-based controllers for swing suppression (modeling_and_observer_based). While these works demonstrate methods for improving crane stability, most rely on precise mechanical actuation, detailed system models, or structural modifications that are impractical for typical mobile crane operations. Prior work has shown that FIR-based filtering can reduce sway during crane motion, but it does not address dynamic stability envelopes for tip-over (digital_filter). Tip-over can be avoided by reducing payload swing through automated trajectory planning when accurate models are available and environmental conditions are fully known. Recent robotic crane systems employ fully autonomous path planning, obstacle avoidance, and motion compensation to achieve reliable point-to-point load transport (robotic_tower_cranes; lifting_path_planning; optimal_traj_planning). These autonomous systems operate without continuous human involvement, relying on pre-planned trajectories. However, fully removing human control overlooks the operator’s intuitive judgment, which is essential in unstructured and dynamic construction environments. Therefore, there is a need for approaches that reduce swing-induced tip-over risk by augmenting human control instead of replacing it. In this work, control techniques are applied to assist manual slewing operations by reducing swing-induced forces while maintaining the operator’s real-time decision authority.
@@
The main contribution of this work is that it mitigates the tip-over of mobile cranes caused by dynamic effects. The method shapes the operator’s motion commands (the input) to cancel vibrations at the system’s natural frequency of the swing dynamics. Specifically, a Multi-Mode Unity-Magnitude Zero-Vibration (MUMZV) input shaper (mumzv) is applied to the slewing velocity command to suppress the swing mode. A major advantage of the proposed approach (using input shaping to reduce payload swing) is that it does not require a detailed crane model. Since the dominant swing frequency depends primarily on the rope length, the input-shaping filter can be designed directly based on this parameter, making it effectively model-free and straightforward to implement. By ensuring that changes in slewing speed do not excite the natural frequencies of the suspended payload, the shaped command minimizes residual oscillations and keeps the motion within safe bounds. Consequently, the MUMZV-based human-in-the-loop approach enables higher slewing speeds while avoiding tip-over. The simulation and experimental results show that the proposed approach substantially recovers the operational-safety envelope of mobile cranes (designed using static analysis) that would otherwise be lost in dynamic conditions.
@@ 3 Methodology
Payload swing behaves as a pendulum that is directly forced by the crane’s slewing acceleration (dynamics_and_control). Applying the Euler–Lagrange equations to the crane–pendulum model gives the swing dynamics (command_shaping_boom_crane) as,
@@ 3.1 Acceleration–Domain MUMZV Input Shaping
Payload swing is driven by the slewing acceleration $\ddot{\alpha}(t)$ input, so input shaping is applied in the acceleration domain to not resonate the swing dynamics. Let $r(t)$ denote the operator’s desired slewing acceleration. A MUMZV shaper (mumzv) is defined by three impulse time instants $t_{1}$, $t_{2}$, $t_{3}$ with corresponding amplitudes $A_{1}$, $A_{2}$, $A_{3}$ that satisfy
@@ 4.2.1 Experimental Setup
A physical experiment was conducted using the scaled mobile crane model shown in Fig. 6. The boom length $L_{b}=0.9144$ m and operating radius $R=0.7$ m were fixed, and the payload mass was set to $m=0.5$ kg. The suspended rope length was $l=0.5715$ m. The crane was placed in an obstacle course designed to represent a realistic operating environment. A video of two cars crossing nearby was used (Pixabay_StopSign_2016) to simulate roadside traffic, and additional obstacles were placed around the crane to create potential collision hazards as shown in Fig. 8. The experimental platform uses the same geometric dimensions as the simulation model. The payload was chosen to be smaller than the statically stable payload to illustrate the potential for tip-over due to swing.

### v2

@@ 1 Introduction
Mobile cranes are widely utilized in construction, shipping, and industrial operations, where both safety and productivity are crucial. Despite their versatility, mobile cranes remain highly susceptible to instability, with tip-over accidents continuing to be a major cause of fatalities and equipment damage (Fujioka et al. (2009); Rauch et al. (2013); Merinar and Moore (2006); Occupational Safety and Health Administration (OSHA) (2025); Abo-Shanab and Sepehri (2006); Rauch (2008)). A significant contributor to such incidents is the dynamic interaction between the boom and the suspended payload during slewing motion. In practice, load limits for crane safety are prescribed by load charts that are based on static equilibrium analysis (Maxim Crane Works (2021); NCCCO (2014)) to specify the maximum allowable load as a function of crane configuration. These limits do not account for dynamic effects such as payload swing due to slewing. As a result, crane operations that are theoretically within the static safe limits may still become unstable due to dynamic effects, as shown in Fig. 1. The goal of this work is to avoid tip-over due to excessive dynamics-induced payload swing during slewing, even with statically safe loads.
@@
Tip-over can be mitigated by modifying the crane’s mechanical configuration or by automating its motion. For example, increasing the counterweight mass has been shown to expand the stability margin and improve lifting capacity in crawler cranes (Zhou et al. (2021)). Other approaches focus on controlling outrigger forces to stabilize the crane during motion (Yang et al. (2023)), though moving outriggers during an operation is generally unsafe. Additional studies examine specific crane architectures and derive model-based controllers for swing suppression (Dŏng and Tuàn (2022)). While these works demonstrate methods for improving crane stability, most rely on precise mechanical actuation, detailed system models, or structural modifications that are impractical for typical mobile crane operations. Prior work has shown that FIR-based filtering can reduce sway during crane motion, but it does not address dynamic stability envelopes for tip-over (Glossiotis and Antoniadis (2007)). Tip-over can be avoided by reducing payload swing through automated trajectory planning when accurate models are available and environmental conditions are fully known. Recent robotic crane systems employ fully autonomous path planning, obstacle avoidance, and motion compensation to achieve reliable point-to-point load transport (Ku et al. (2024); Zhou et al. (2021); Chen et al. (2016)). These autonomous systems operate without continuous human involvement, relying on pre-planned trajectories. However, fully removing human control overlooks the operator’s intuitive judgment, which is essential in unstructured and dynamic construction environments. Therefore, there is a need for approaches that reduce swing-induced tip-over risk by augmenting human control instead of replacing it. In this work, control techniques are applied to assist manual slewing operations by reducing swing-induced forces while maintaining the operator’s real-time decision authority.
@@
The main contribution of this work is that it mitigates the tip-over of mobile cranes caused by dynamic effects. The method shapes the operator’s motion commands (the input) to cancel vibrations at the system’s natural frequency of the swing dynamics. Specifically, a Modified Unity-Magnitude Zero-Vibration (MUMZV) input shaper (Sung and Singhose (2008)) is applied to the slewing velocity command to suppress the swing mode. A major advantage of the proposed approach (using input shaping to reduce payload swing) is that it does not require a detailed crane model. Since the dominant swing frequency depends primarily on the rope length, the input-shaping filter can be designed directly based on this parameter, making it effectively model-free and straightforward to implement. By ensuring that changes in slewing speed do not excite the natural frequencies of the suspended payload, the shaped command minimizes residual oscillations and keeps the motion within safe bounds. Consequently, the MUMZV-based human-in-the-loop approach enables higher slewing speeds while avoiding tip-over. The simulation and experimental results show that the proposed approach substantially recovers the operational-safety envelope of mobile cranes (designed using static analysis) that would otherwise be lost in dynamic conditions.
@@ 3 Methodology
Payload swing behaves as a pendulum that is directly forced by the crane’s slewing acceleration (Abdel-Rahman et al. (2003)). Applying the Euler–Lagrange equations to the crane–pendulum model gives the swing dynamics (Parker et al. (1999)) as,
@@ 3.1 Acceleration–Domain MUMZV Input Shaping
Payload swing is driven by the slewing acceleration $\ddot{\alpha}(t)$ input, so input shaping is applied in the acceleration domain to not resonate the swing dynamics. Let $r(t)$ denote the operator’s desired slewing acceleration. A MUMZV shaper (Sung and Singhose (2008)) is defined by three impulse time instants $t_{1}$, $t_{2}$, $t_{3}$ with corresponding amplitudes $A_{1}$, $A_{2}$, $A_{3}$ that satisfy
@@ 4.2.1 Experimental Setup
A physical experiment was conducted using the scaled mobile crane model shown in Fig. 6. The boom length $L_{b}=0.9144$ m and operating radius $R=0.7$ m were fixed, and the payload mass was set to $m=0.5$ kg. The suspended rope length was $l=0.5715$ m. The crane was placed in an obstacle course designed to represent a realistic operating environment. A video of two cars crossing nearby was used (AzamKamolov (2016)) to simulate roadside traffic, and additional obstacles were placed around the crane to create potential collision hazards as shown in Fig. 8. The experimental platform uses the same geometric dimensions as the simulation model. The payload was chosen to be smaller than the statically stable payload to illustrate the potential for tip-over due to swing.